# Compression fronts from fast radio bursts

### v1

@@ 4 MHD equations for fluid motion
In the test-particle numerical experiments described above, we prescribed an arbitrary pattern of the fluid drift speed $\beta_{\rm D}(\xi)$. Our goal now is to find the actual pattern of fluid motion in the radio wave packet. It can be quantified by the profiles of two quantities: $\kappa$ (related to $\beta_{\rm D}$) and $\tilde{C}=\tilde{\rho}/\tilde{\rho}_{\rm u}$ (the proper compression factor). They are equal, $\tilde{C}=\kappa$, when a steady pattern of fluid motion is established in the propagating wave packet, i.e. when $\kappa$ and $\tilde{\rho}$ are functions of $\xi$ only (see below). The full problem, however, involves evolving profiles. Therefore, we will formulate general MHD equations for $\kappa(t,\xi)$ and $\tilde{\rho}(t,\xi)$, allowing time dependence. We first state the equations in the usual variables $t,z$ and later switch to the more convenient coordinates $t,\xi$.
4.1 Conservation laws
@@
The relevant components of $T^{\mu\nu}$ are $T^{tt}$, $T^{tz}$, and $T^{zz}$. In particular, the wave stress-energy tensor is
@@
The plasma stress-energy tensor is
$T^{\mu\nu}_{\rm p}=(\tilde{\rho}c^{2}+U_{\rm p}+P_{\rm p})u_{\rm D}^{\mu}u_{\rm D}^{\nu}+\eta^{\mu\nu}P_{\rm p},$ (26)
where $\eta^{\mu\nu}$ is the Minkowski metric. The plasma pressure $P_{\rm p}$ and internal energy $U_{\rm p}$ (Equation 22) are controlled by the wave strength $a(\xi)$. In particular, $\tilde{\gamma}=\sqrt{1+a^{2}}$ in the regular heating regime ($\tilde{\omega}_{B}<\tilde{\omega}$).
Relevant components of the stress-energy tensor of the MHD electromagnetic field are
$T^{tt}_{\rm f}=T^{zz}_{\rm f}=\frac{B^{2}+E^{2}}{8\pi},\qquad T^{tz}_{\rm f}=\frac{EB}{4\pi},$ (27)
@@
$\frac{\tilde{B}}{B_{\rm u}}=\frac{\tilde{\rho}}{\rho_{\rm u}}.$ (30)
The total MHD stress-energy tensor may be summarized as follows
$T^{\mu\nu}_{\rm MHD}=T^{\mu\nu}_{\rm p}+T^{\mu\nu}_{\rm f}=h\tilde{\rho}c^{2}u_{\rm D}^{\mu}u_{\rm D}^{\nu}+\eta^{\mu\nu}P,$ (31)
where
$h=1+w+\sigma,\qquad P=\tilde{\rho}c^{2}\left(\frac{w}{3}+\frac{\sigma}{2}\right),$ (32)
$w\equiv\frac{U_{\rm p}+P_{\rm p}}{\tilde{\rho}c^{2}}=\frac{3}{2}(\tilde{\gamma}-1),\quad\sigma\equiv\frac{\tilde{B}^{2}}{4\pi\tilde{\rho}c^{2}}=\frac{\tilde{\rho}}{\tilde{\rho}_{\rm u}}\,\sigma_{\rm u}.$ (33)
Note that the stress-energy tensor of the MHD fluid depends on the strength parameter $a$ of the wave it interacts with. In particular, in the regular heating regime, Equation (9) implies that the dimensionless enthalpy $w$ tracks the local strength of the wave $a(\xi)$:
$w\approx\frac{3}{2}a\qquad\left({\rm when~{}}\,\tilde{\omega}_{B}<\tilde{\omega}~{}{\rm and}~{}a\gg 1\right).$ (34)
@@
$\partial_{\nu}\left(T^{\mu\nu}_{\rm MHD}+T^{\mu\nu}_{\rm w}\right)=-\tilde{n}\dot{{\cal E}}_{e}u_{\rm D}^{\mu}.$ (35)
The term $\tilde{n}\dot{{\cal E}}_{e}u^{\mu}$ is the radiative loss rate (per unit volume) of the plasma particles, which oscillate and gyrate in the wave; $\tilde{n}=\tilde{\rho}/m$ is the particle number density and ˙$\dot{{\cal E}}_{e}$ is the emitted power per particle, which is a Lorentz-invariant quantity.
The power ˙$\dot{{\cal E}}_{e}$ can be calculated in the fluid frame $\tilde{\cal K}$. When $\tilde{\omega}_{B}<\tilde{\omega}$, it is given by
$\dot{\gamma}_{e}\equiv\frac{\dot{{\cal E}}_{e}}{mc^{2}}\approx\frac{2}{3}\frac{r_{e}}{c}\tilde{\omega}^{2}a^{2}\tilde{\gamma}^{2}.$ (36)
A similar expression (with the numerical coefficient $4/3$ instead of $2/3$) holds in the regime of stochastic heating, $\tilde{\omega}_{B}>\tilde{\omega}$ (B22). The strength parameter of the radio wave $a$ is invariant, and its frequency $\omega$ and electric field $E_{\rm wave}$ transform to the fluid frame as
$\frac{\tilde{\omega}}{\omega}=\frac{\tilde{E}_{\rm wave}}{E_{\rm wave}}=\kappa^{-1},$ (37)
where $\kappa$ is defined in Equation (13). Equation (36) may also be written in terms of the wave energy density $U$,
$\dot{\gamma}_{e}\approx\frac{\sigma_{\rm T}\tilde{U}}{mc}\,\tilde{\gamma}^{2}=\frac{\sigma_{\rm T}U}{mc\,\kappa^{2}}\,\tilde{\gamma}^{2},$ (38)
where $\sigma_{\rm T}=(8\pi/3)r_{e}^{2}$ is the Thomson cross section.
4.2 MHD equations in coordinates $t,\xi$
@@
Similarly, the flux of four-momentum ${\cal P}^{\mu}$ through a surface $\xi=const$ is $T^{\mu t}-T^{\mu z}$. Conservation of four-momentum (Equation 35) in the $t,\xi$-coordinates becomes [^6] This equation can also be derived formally by transforming tensor $T^{\mu\nu}$ to new coordinates: $(t,z)\rightarrow(t,\xi)$. The resulting $T^{t\xi}$ and $T^{z\xi}$ are the energy and momentum fluxes crossing $\xi=const$.
$\left.\frac{\partial T^{\mu t}}{\partial t}\right|_{\xi}+\frac{\partial}{\partial\xi}\left(T^{\mu t}-T^{\mu z}\right)_{t}=-\tilde{n}\dot{{\cal E}}_{e}u_{\rm D}^{\mu}.$ (41)
Note that the wave is static in $\xi$, so it creates no flux of energy or momentum across a surface of $\xi=const$:
$T^{\mu t}_{\rm w}-T^{\mu z}_{\rm w}=U-U=0\qquad(\mu=t,z).$ (42)
Only the MHD fluid contributes to the fluxes. Thus, the fluxes of energy and momentum through a surface of $\xi=const$ are
$\frac{F_{\rm en}}{c}$ $=T^{tt}_{\rm MHD}-T^{tz}_{\rm MHD}=\frac{c^{2}\rho h}{\kappa}-P,$ (43)
$F_{\rm mom}$ $=T^{zt}_{\rm MHD}-T^{zz}_{\rm MHD}=\frac{c^{2}\rho h}{\kappa}\beta_{\rm D}-P.$ (44)
4.3 Conservation of ${\cal Q}_{\pm}$
Instead of energy ${\cal E}$ and momentum ${\cal P}$ it will be helpful to use quantities
${\cal Q}_{\pm}={\cal E}\pm c{\cal P}.$ (45)
The fluxes of ${\cal Q}_{\pm}$ through a surface of $\xi=const$ are
$F_{-}$ $=F_{\rm en}-cF_{\rm mom}=\frac{c^{3}\tilde{\rho}h}{\kappa^{2}},$ (46)
$F_{+}$ $=F_{\rm en}+cF_{\rm mom}=c^{3}\tilde{\rho}h-2cP.$ (47)
@@ 4.3 Conservation of ${\cal Q}_{\pm}$
Conservation of ${\cal Q}_{\pm}$ is given by the sum/difference of equations describing energy and momentum conservation (Equation (41) with $\mu=t$ and $z$):
$\left.\frac{\partial}{\partial t}\left(T^{tt}\pm T^{tz}\right)\right|_{\xi}+\frac{\partial}{\partial\xi}\left(\frac{F_{\pm}}{c}\right)_{t}=-\tilde{n}\dot{{\cal E}}_{e}(u_{\rm D}^{t}\pm u_{\rm D}^{z}).$ (48)
Note that the stress-energy tensor of the radio wave disappears from the conservation law for ${\cal Q}_{-}$. This has a simple physical reason: the energy and momentum of the wave satisfy ${\cal E}_{\rm w}=c{\cal P}_{\rm w}$, so the wave carries ${\cal Q}_{-}=0$. There is no exchange of ${\cal Q}_{-}$ in the wave-fluid interaction: whenever the radio wave loses energy $\Delta{\cal E}$ to the MHD fluid, it also loses momentum $\Delta{\cal P}=\Delta{\cal E}/c$, so
$\Delta{\cal Q}_{-}=\Delta{\cal E}-c\,\Delta{\cal P}=0.$ (49)
@@
This observation significantly simplifies the problem. As long as the shape of the propagating radio wave $a(\xi)$ is approximately steady (i.e. the wave is not damped), the fluid dynamics in the packet is fully described by two equations that express conservation of mass and ${\cal Q}_{-}$:
@@
$\partial_{t}\left(\frac{\rho h}{\kappa}-\frac{P}{c^{2}}\right)$ $=-\partial_{\xi}\left(\frac{\tilde{\rho}h}{\kappa^{2}}\right)-\frac{\tilde{\rho}}{\kappa}\dot{\gamma}_{e}.$ (51)
@@
The radio wave does not contribute to these conservation laws, because it carries no mass or ${\cal Q}_{-}$. The wave strength parameter $a(\xi)$ still enters the ${\cal Q}_{-}$ equation through $h=1+w+\sigma$ (since $a$ controls the fluid enthalpy $w$) and through $\dot{\gamma}_{e}\propto a^{2}$. As long as $a(\xi)$ is steady, it can be treated as a given function in the wave-fluid interaction problem, so Equations (50) and (51) form a closed set for two unknowns $\rho(t,\xi)$ and $\kappa(t,\xi)$. For a given radio wave packet, the equations can be solved in the domain $0<\xi<T$ with given initial conditions $\rho(0,\xi)$ and $\kappa(0,\xi)$, and boundary conditions $\rho(t,0)=\rho_{\rm u}$ and $\kappa(t,0)=\kappa_{\rm u}$.
Equations (50) and (51) allow one to focus on the fluid dynamics in the wave packet without considering the evolution of the radio wave itself, assuming that the evolution of $a(\xi)$ is negligible on timescales of interest. The consistency of this assumption can be checked later using the remaining independent equation that states conservation of ${\cal Q}_{+}$ and governs the evolution of $a(t,\xi)$. A simpler consistency check can be done by comparing the energy exchanged in the wave-plasma interaction with the wave energy (Section 10).
@@
Note also that Equations (50) and (51) are Lorentz-invariant: they have the same form in any frame boosted along $z$ [^7] Lorentz transformation of initial conditions is not, however, straightforward, since hypersurfaces of simultaneity ($t=const$) are different in different frames. They can be solved e.g. in the frame where the upstream plasma is at rest: $\kappa_{\rm u}=1$ and $\rho_{\rm u}=\tilde{\rho}_{\rm u}$. The problem can also be solved in other frames where $\kappa_{\rm u}\neq 1$ and the upstream plasma has speed $\beta_{\rm u}=(\kappa_{\rm u}^{2}-1)/(\kappa_{\rm u}^{2}+1)$, Lorentz factor $\gamma_{\rm u}=(\kappa_{\rm u}^{2}+1)/2\kappa_{\rm u}$, and proper density $\tilde{\rho}_{\rm u}=\rho_{\rm u}/\gamma_{\rm u}$.
5 Steady compression front
@@
The pattern of fluid motion inside the propagating wave packet will be called below the “front” for brevity. Its structure is described by the profiles of $\rho(t,\xi)$ and $\kappa(\xi,t)$, which are governed by Equations (50) and (51). In general, the front structure can evolve in time, as demonstrated in the next section. Here, we first show that there is a unique steady flow pattern, where all quantities depend on $\xi$ only, with no evolution with $t$. As verified in section 6, the steady state is an attractor for the front evolution. It is reached under two conditions: (1) the front propagation time is long enough for the flow pattern to relax to the steady state, and (2) the propagation time is short enough to neglect damping of the wave, so $a(\xi)$ can be treated as constant in time.
@@
$\frac{d}{d\xi}\left(\frac{h}{\kappa}\right)$ $=-\dot{\gamma}_{e}.$ (53)
@@
The solution $\tilde{\rho}/\rho_{\rm u}=\kappa/\kappa_{\rm u}$ also implies $\sigma=\sigma_{\rm u}\kappa/\kappa_{\rm u}$ and $d/d\xi(h/\kappa)=d/d\xi[(1+w)/\kappa]$. Thus, Equation (53) becomes
$\frac{d}{d\xi}\left(\frac{1+w}{\kappa}\right)=-\dot{\gamma}_{e}=-\frac{\sigma_{\rm T}U}{mc}\,\frac{\tilde{\gamma}^{2}}{\kappa^{2}}.$ (56)
Note that the fluid magnetization $\sigma$ has dropped out. Remarkably, the fluid motion and compression in a steady front are independent of $\sigma_{\rm u}$. In particular, in the case of negligible radiative losses, we find
$\kappa=(1+w)\kappa_{\rm u}\qquad({\rm when}~{}\,\dot{\gamma}_{e}=0),$ (57)
where we used the cold upstream condition $w_{\rm u}=0$.
One can also solve for $\kappa(\xi)$ without neglecting radiative losses. In particular, consider the simpler regime of regular heating ($\tilde{\omega}_{B}<\tilde{\omega}_{B}$). We are interested in waves with $a\gg 1$, which heat the plasma to $w\approx(3/2)\tilde{\gamma}\gg 1$. Equation (56) can be integrated for $(1+w)/\kappa$, and we find
$\kappa=(1+w)\kappa_{\rm u}(1+q),$ (58)
@@
[…]
@@
The importance of radiative losses is controlled by the (Lorentz-invariant) dimensionless parameter $q$, which is determined by the column energy density of the wave packet, $\Sigma$. Equation (57) holds when $q\ll 1$.
6 Relaxation to the steady state

### v2

@@ IV MHD equations for fluid motion
In the test-particle numerical experiments described above, we prescribed an arbitrary pattern of the fluid drift speed $\beta_{\rm D}(\xi)$ or, equivalently, $\kappa(\xi)$. Our goal now is to find the actual pattern of fluid motion in the radio wave packet. It can be quantified by the profiles of two quantities: $\kappa$ (related to $\beta_{\rm D}$) and $\tilde{C}=\tilde{\rho}/\tilde{\rho}_{\rm u}$ (the proper compression factor). They turn out equal, $\tilde{C}=\kappa$, when a steady pattern of fluid motion is established in the propagating wave packet, i.e. when $\kappa$ and $\tilde{\rho}$ are functions of $\xi$ only. The full problem, however, involves evolving profiles. Therefore, we will formulate general MHD equations for $\kappa(t,\xi)$ and $\tilde{\rho}(t,\xi)$, allowing time dependence. We first state the equations in the usual variables $t,z$ and later switch to the more convenient coordinates $t,\xi$.
IV.1 Conservation laws
@@
The relevant components of $T^{\mu\nu}$ are $T^{tt}$, $T^{tz}$, and $T^{zz}$. In particular, the stress-energy tensor of the radio wave is
@@
Consider now the relevant components of the stress-energy tensor of the MHD electromagnetic field:
$T^{tt}_{\rm f}=T^{zz}_{\rm f}=\frac{B^{2}+E^{2}}{8\pi},\qquad T^{tz}_{\rm f}=\frac{EB}{4\pi}.$ (30)
@@
$\frac{\tilde{B}}{B_{\rm u}}=\frac{\tilde{\rho}}{\rho_{\rm u}}=\frac{\sigma}{\sigma_{\rm u}},\qquad\sigma\equiv\frac{\tilde{B}^{2}}{4\pi\tilde{\rho}c^{2}}.$ (33)
The remaining component of the system is the plasma, which is energized by the radio wave. The plasma stress-energy tensor $T_{\rm p}^{\mu\nu}$ is derived in Appendix C for both possible regimes: stochastic heating and regular oscillation in the wave (Equations (C6) and (C8), respectively). The total MHD stress-energy tensor is
$T^{\mu\nu}_{\rm MHD}=T^{\mu\nu}_{\rm p}+T^{\mu\nu}_{\rm f}.$ (34)
@@
$\partial_{\nu}\left(T^{\mu\nu}_{\rm MHD}+T^{\mu\nu}_{\rm w}\right)=-n\dot{{\cal P}}_{e}^{\mu},$ (35)
where $n=\rho/m$ is the particle number density, and $\dot{{\cal P}}_{e}^{\mu}=(\dot{{\cal E}}_{e}/c,\dot{\boldsymbol{{\cal P}}}_{e})$ is the average emission rate of four-momentum per particle, which is derived in Appendix D. It is given by Equations (D9), (D10) in the regular oscillation regime and by Equations (D19), (D20) in the stochastic heating regime.
IV.2 MHD equations in coordinates $t,\xi$
@@
Similarly, the flux of four-momentum ${\cal P}^{\mu}$ through a surface $\xi=const$ is $F^{\mu}=T^{\mu t}-T^{\mu z}$. Conservation of four-momentum (Equation 35) in the $t,\xi$-coordinates becomes [^7] This equation can also be derived formally by transforming tensor $T^{\mu\nu}$ to new coordinates: $(t,z)\rightarrow(t,\xi)$. The resulting $T^{t\xi}$ and $T^{z\xi}$ are the energy and momentum fluxes crossing $\xi=const$.
$\left.\frac{\partial T^{\mu t}}{\partial t}\right|_{\xi}+\frac{\partial}{\partial\xi}\left(T^{\mu t}-T^{\mu z}\right)_{t}=-cn\dot{{\cal P}}_{e}^{\mu}.$ (38)
Note that the wave is static in $\xi$, so it creates no flux of energy or momentum across a surface of $\xi=const$: [^8] This statement holds for waves propagating with speed $c$. The exact propagation speed $v_{\rm gr}<c$ (Appendix A) implies that the radio wave carries a non-zero energy flux in $\xi$: $F_{\rm w}=U(c-v_{\rm gr})$. In particular, in the fluid frame $\tilde{\cal K}$, $\tilde{F}_{\rm w}\sim c\tilde{U}\tilde{\omega}_{\rm p}^{2}/2\tilde{\omega}^{2}\sim\tilde{\rho}c^{3}a^{2}/2\tilde{\gamma}$. One may neglect $\tilde{F}_{\rm w}$ compared to flux $\tilde{F}$ created by the MHD fluid (effectively setting $v_{\rm gr}=c$) if $a^{2}/\tilde{\gamma}\ll h$, where $h$ is given in Equation (44) below. This condition turns out satisfied if the plasma has $\sigma_{\rm u}\gg 1$.
$F^{\mu}_{\rm w}=T^{\mu t}_{\rm w}-T^{\mu z}_{\rm w}=U-U=0\qquad(\mu=t,z).$ (39)
Then, only the MHD fluid contributes to the fluxes.
IV.3 Conservation of $\cal Q$$^{\pm}$
[…]
${\cal Q}^{\pm}={\cal E}\pm c{\cal P}.$ (40)
@@ IV.3 Conservation of $\cal Q$$^{\pm}$
Conservation of $\cal Q$$^{\pm}$ is given by the sum/difference of equations stating energy and momentum conservation (Equation (38) with $\mu=t$ and $z$):
$\frac{\partial}{\partial t}(T^{tt}+T^{tz})+\frac{\partial}{\partial\xi}(T^{tt}-T^{zz})$ $=-\dot{\cal Q}^{+},$ (41)
$\frac{\partial}{\partial t}(T^{tt}-T^{tz})+\frac{\partial}{\partial\xi}(T^{tt}-2T^{tz}+T^{zz})$ $=-\dot{\cal Q}^{-},$ (42)
where $\dot{\cal Q}^{\pm}=cn(\dot{{\cal P}}_{e}^{t}\pm\dot{{\cal P}}_{e}^{z})$.
[…]
$\Delta{\cal Q}^{-}=\Delta{\cal E}-c\,\Delta{\cal P}=0.$ (43)
@@
This observation significantly simplifies the problem. As long as the propagating radio wave packet $a(\xi)$ is approximately steady (i.e. the wave is not damped), the fluid dynamics in the packet is fully described by two equations that express conservation of mass and $\cal Q$$^{-}$. The remaining independent equation for $Q^{+}$ is needed if one aims to explicitly follow the gradual evolution of the wave packet $a(t,\xi)$. It is not needed for FRBs that are weakly affected by their interaction with the wind. Here, we will assume this condition to be satisfied, and later (in section X) will estimate the parameter space where strong FRB damping is expected.
Another observation concerns the form of the plasma stress-energy tensor $T_{\rm p}^{\mu\nu}$ derived in Appendix C. In the stochastic heating regime, it takes the usual ideal-fluid form (Equation C8). By contrast, in the regular oscillation regime, $T_{\rm p}^{\mu\nu}$ looks different because of the first term in Equation (C6). However, this term is the same in all three components $T_{\rm p}^{tt}$, $T_{\rm p}^{tz}$, $T_{\rm p}^{zz}$ and drops out from Equation (42). Thus, effectively, only the second term in $T_{\rm p}^{\mu\nu}$ contributes to the $Q^{-}$ equation, i.e. the plasma stress-energy tensor can be replaced by $T_{\rm p}^{\mu\nu}=(1+a^{2})^{1/2}\tilde{\rho}c^{2}u_{\rm D}^{\mu}u_{\rm D}^{\nu}$. It has the ideal-fluid form with zero pressure and non-zero dimensionless enthalpy $w$ defined by $1+w=(1+a^{2})^{1/2}$. Recall that the MHD electromagnetic field also has $T_{\rm f}^{\mu\nu}$ of an ideal-fluid form (Equation 31), with effective pressure $\tilde{\rho}c^{2}\sigma/2$ and enthalpy $\tilde{\rho}c^{2}\sigma$.
As a result, one can replace $T^{\mu\nu}$ in Equation (42) with the effective $T_{\rm MHD}^{\mu\nu}=\tilde{\rho}c^{2}hu_{\rm D}^{\mu}u_{\rm D}^{\nu}+P\eta^{\mu\nu}$ with the following definitions of $h$ and $P$:
$h=1+w+\sigma,$ (44)
$w=\sqrt{1+a^{2}}-1,\qquad\frac{P}{\tilde{\rho}c^{2}}=\frac{\sigma}{2}\qquad({\rm regular})$ (45a)
$w=\frac{3}{2}(\tilde{\gamma}-1),\qquad\frac{P}{\tilde{\rho}c^{2}}=\frac{\sigma}{2}+\frac{w}{3}\quad\;({\rm stochastic})$ (45b)
In the regular oscillation regime, the plasma makes no contribution to pressure $P$; the oscillations only increase the effective inertial mass of each particle by the factor $\tilde{\gamma}=\sqrt{1+a^{2}}$. Unlike stochastic heating, coherent oscillations are not true heating. However, it is convenient to define effective enthalpy $w$ in both regimes, so that they have the common form of $T_{\rm MHD}^{\mu\nu}$.
The equations of mass and $Q^{-}$ conservation can now be stated as follows
@@
$\partial_{t}\left(\frac{\rho h}{\kappa}-\frac{P}{c^{2}}\right)$ $=-\partial_{\xi}\left(\frac{\tilde{\rho}h}{\kappa^{2}}\right)-\frac{n\dot{{\cal Q}}^{-}_{e}}{c^{2}}.$ (47)
The term $\dot{{\cal Q}}^{-}_{e}=c(\dot{{\cal P}}_{e}^{t}-\dot{{\cal P}}_{e}^{z})$ describes radiative losses per particle. It is determined by Equations (D9), (D10) and Equations (D19), (D20) in the regimes of regular oscillation and stochastic heating:
$\!\!\!\!\!\!\dot{\cal Q}^{-}_{e}\!=\!\frac{c\sigma_{\rm T}U}{\gamma_{\rm D}\kappa^{3}}\!\!\times\!\!\left\{\begin{array}[]{lr}\vskip 2.84526pt1+a^{2}&\!\!\!\!\!{\rm regular}\\ \displaystyle{\frac{5}{2}\tilde{\gamma}^{2}}&\!\!\!\!{\rm stochastic}\end{array}\!\right\}\!\approx\!\frac{c\sigma_{\rm T}U}{\gamma_{\rm D}\kappa^{3}}(1+w)^{2}.$ (50)
where $\sigma_{\rm T}=(8\pi/3)r_{e}^{2}$ is the Thomson cross section.
@@
The radio wave does not contribute to Equations (46) and (47), because it carries no mass and negligible $\cal Q$$^{-}$. The wave strength parameter $a(\xi)$ still enters the $\cal Q$$^{-}$ equation through $h=1+w+\sigma$ (since $a$ controls the fluid enthalpy $w$) and through $\dot{{\cal Q}}_{e}^{-}\propto U\propto a^{2}$. As long as $a(\xi)$ is steady (not damped), it can be treated as a given function in the wave-fluid interaction problem, and Equations (46) and (47) form a closed set for two unknowns $\rho(t,\xi)$ and $\kappa(t,\xi)$. The equations can be solved in the domain $0<\xi<T$ (i.e. inside the radio wave packet) with given initial conditions $\rho(0,\xi)$ and $\kappa(0,\xi)$, and boundary conditions $\rho(t,0)=\rho_{\rm u}$ and $\kappa(t,0)=\kappa_{\rm u}$.
@@
Equations (46) and (47) are Lorentz-invariant: they have the same form in any frame boosted along zz. [^9] Lorentz transformation of initial conditions is not, however, trivial, since hypersurfaces of simultaneity ($t=const$) are different in different frames. They can be solved e.g. in the frame where the upstream plasma is at rest: $\kappa_{\rm u}=1$ and $\rho_{\rm u}=\tilde{\rho}_{\rm u}$. The problem can also be solved in other frames where $\kappa_{\rm u}\neq 1$ and the upstream plasma has speed $\beta_{\rm u}=(\kappa_{\rm u}^{2}-1)/(\kappa_{\rm u}^{2}+1)$, Lorentz factor $\gamma_{\rm u}=(\kappa_{\rm u}^{2}+1)/2\kappa_{\rm u}$, and density $\rho_{\rm u}=\gamma_{\rm u}\tilde{\rho}_{\rm u}$.
V Steady compression front
@@
The pattern of fluid motion inside the propagating wave packet will be called below the “front” for brevity. Its structure is described by the profiles of $\rho(t,\xi)$ and $\kappa(\xi,t)$, which are governed by Equations (46) and (47). In general, the front structure can evolve in time, as demonstrated in the next section. Here, we first show that there is a unique steady flow pattern, where all quantities depend on $\xi$ only, with no evolution with $t$. As verified in section VI, the steady state is an attractor for the front evolution. It is reached under two conditions: (1) the front propagation time is long enough for the flow pattern to relax to the steady state, and (2) the propagation time is short enough to neglect damping of the radio wave, so $a(\xi)$ can be treated as constant in time.
@@
$\frac{d}{d\xi}\left(\frac{h}{\kappa}\right)$ $=-\frac{\kappa\gamma_{\rm D}\dot{{\cal Q}}^{-}_{e}}{mc^{2}}.$ (52)
@@
The solution $\tilde{\rho}/\tilde{\rho}_{\rm u}=\kappa/\kappa_{\rm u}$ also implies $\sigma=\sigma_{\rm u}\kappa/\kappa_{\rm u}$ and $d/d\xi(h/\kappa)=d/d\xi[(1+w)/\kappa]$. As a result, Equation (52) becomes
$\frac{d}{d\xi}\left(\frac{1+w}{\kappa}\right)=-\frac{\sigma_{\rm T}U(1+w)^{2}}{mc\,\kappa^{2}}.$ (55)
Note that the fluid magnetization $\sigma$ has dropped out. Remarkably, fluid motion and compression in a steady front are unaffected by the magnetic field, regardless of how strong it is. Integrating Equation (55) for $(1+w)/\kappa$ and using the cold upstream condition $w_{\rm u}=0$, we find
$\kappa=(1+q)(1+w)\kappa_{\rm u},$ (56)
@@
$q(\xi)=\frac{\sigma_{\rm T}\Sigma(\xi)}{mc^{2}\kappa_{\rm u}},\qquad\Sigma(\xi)\equiv\int_{0}^{\xi}U(\xi^{\prime})c\,d\xi^{\prime}.$ (57)
@@
The effect of radiative losses on the fluid motion is controlled by the (Lorentz-invariant) dimensionless parameter $q$, which is determined by the column energy density of the radio wave packet, $\Sigma$. Losses weakly affect the front when $q\ll 1$.
In the regular oscillation regime with $q\ll 1$, the obtained steady-state solution for the compression front simplifies to
$\kappa=\kappa_{\rm u}\sqrt{1+a^{2}}\qquad\mathrm{(regular\;\&\;non\mathrm{-}radiative)}.$ (58)
In this limit, the solution agrees with the result of Sobacchi et al. (2024). This agreement is remarkable, since Sobacchi et al. (2024) used a different approach to the steady-state compression front: they studied plasma in the frame moving with the group speed of a nearly monochromatic wave packet, and employed an expansion in Fourier harmonics, assuming that the upstream plasma ahead of the packet has $\tilde{\omega}_{B}<\tilde{\omega}/a$.
VI Relaxation to the steady state